OpenStudy (anonymous):

I WILL GIVE YOU A MEDAL AND FAN YOU! Choose the correct simplification of the expression (x2y)2. x4y3 x6y6 x4y2 xy4

OpenStudy (mathstudent55):

You need these two rules: Rule: \( (a^m)^n = a^{mn} \) Rule: \((ab)^m = a^mb^m \)

OpenStudy (anonymous):

And how do I use those rules @mathstudent55

OpenStudy (johnweldon1993):

Well what those rules state 1) when 2 exponents are separated by parenthesis...they are multiplied 2) when 2 different terms are in the parenthesis...each is raised to that power on the outside of the parenthesis so \[(x^2y)^2\] becomes \[x^{2\times 2} \times y^2 \] so what is x^(2*2)

OpenStudy (anonymous):

x^4y^2

OpenStudy (johnweldon1993):

That would be correct!